# A source of heralded atom-photon entanglement for quantum networking

### v1

@@ Abstract
Our experiment starts by optically cooling and trapping individual atoms at the center of two crossed optical fiber cavities [20]. The “qubit cavity” couples the upper transition and the “herald cavity” couples the lower transition (see Fig. 1). After optically pumping the atom into state , two laser pulses induce a two-photon excitation to state , and the atom subsequently decays, emitting single photons into both cavities [31]. Both cavities have one of the mirrors with a significantly higher transmission than the other one (340 ppm versus 10 ppm), such that photons exit the cavities dominantly through this mirror. During the experiment, we introduce a magnetic field pointing along the qubit cavity, such that the quantization axis is defined in that direction, and single photons in a superposition of right circular and left circular polarizations are emitted into this cavity. The mirrors of the herald cavity are elliptically curved such that the cavity is birefringent with nondegenerate polarization eigenmodes [32]. The direction of the oscillating electric field of one polarization eigenmode (the mode) is oriented along the magnetic field so that this mode enhances the emission of -polarized photons. This mode is resonant to the lower transition while the other herald cavity eigenmode (the V mode) is 500 MHz blue detuned from this transition. In this situation, the two cavities enhance the emission of a single-photon polarization qubit in the qubit cavity that is entangled with the atomic spin and correlated with a -polarized single photon (herald photon) emitted in the herald cavity. The entangled state can be expressed as
@@
In order to characterize the atom-photon entangled state and the impact of the herald photon, Fig. 3 shows the density matrix of the two-qubit state for three different situations. In (a), the herald cavity is not resonant to the lower transition, and all the events with a qubit photon detection are considered. In (b), the herald cavity is resonant to the lower transition, and all the events with a qubit photon detection are considered. In (c) the herald cavity is resonant with the lower transition, too, but only the events with a photon detection in both the qubit and the herald cavities are considered. We observe how the fidelity of the atom-photon entangled state compared to an ideal Bell state increases from in (a) to in (b), and to in (c), well above the fidelity bound for entangled states () [35]. This increase is due to the enhancement of atomic decay on the lower transition, a process which is induced by the mode of the herald cavity, and which can be heralded by the detection of a polarized herald photon. The fidelity difference between the cavity enhanced heralded and unheralded entangled states is due to the finite Purcell factor in the herald cavity transition (=2.66 as measured in [31]). The obtained heralded entangled state fidelity is limited by the decoherence of the atomic state during the atomic qubit measurement time. During this time, the qubit experiences dephasing due to both real magnetic field fluctuations and virtual magnetic field fluctuations induced by the laser traps, as well as differential light shift fluctuations caused by variations in the trap intensity.
@@
In conclusion, we have shown the generation of a heralded atom-photon entangled state using a single atom coupled to two optical cavities. We have observed that heralding enhances the entangled-state generation efficiency and fidelity when compared to the unheralded case. This has the potential to enhance long distance atom-atom entanglement distribution rates by a factor up to . Moreover, we have used the temporal correlation information between the qubit and herald photons in order to reduce the duration of the qubit measurement window. Making use of this feature, we have observed an improvement of the entangled state fidelity from to in a noise-limited photon-qubit measurement situation, which is relevant in long-distance quantum communication.
The achieved fidelity could be improved by having a faster atomic read-out sequence, which could be realized with a more powerful microwave drive or by using laser-driven Raman transitions. A further approach to minimize the decoherence of the atomic qubit is to implement ground state cooling and to employ laser traps with higher detuning from the atomic transitions. In our current system, we must switch off the 852 nm dipole trap during atomic excitation, since the absorption of a trap photon can photoionize the state. This increases the atom’s temperature, enhancing the qubit dephasing and, thereby, reducing the entangled state fidelity. Keeping the trap on would avoid these effects, which would be possible if the transition were employed for the photonic qubit, as the binding energy of the state is higher than a trap photon’s energy.
@@ III Appendix C: Entangled state fidelity in the presence of noise
where is the fidelity of with respect to a Bell state. depends on the probability to detect a noise photon and the probability to detect a qubit photon at the end of the communication line. Due to fiber losses the qubit photon detection probability decreases exponentially with the distance , and therefore the entangled state fidelity also decreases. We can observe this in Fig. 8. The three different curves simulate the distribution of our unheralded atom-photon entangled state, our heralded state, and our heralded state when a herald-adjusted shorter measurement gate time is used (as described in the main text). The three different curves consider the source parameters shown in Fig.3 in the main text, a detection efficiency of and a noise detection rate of 10Hz. The two heralded entanglement fidelity curves consider a photon qubit emission efficiency , and for the unheralded curve . The two heralded entanglement curves are compared to experimental data in Fig. 4c in the main text.
IV Appendix D: State-of-the-art comparison
Table 1 shows a comparison between our work and other state-of-the-art systems capable of generating entanglement between a photon and a quantum emitter/memory. The signal-to-noise ratio (SNR) is affected by both the entanglement generation efficiency and the duration of the photon. In particular, a higher entanglement generation efficiency will lead to a higher SNR, whereas a longer photon duration will set a longer detection window, increasing the likelihood of a dark count and thereby leading to a lower SNR. In order to compare the impact of these two factors across different platforms, we calculate the mean efficiency per time bin. As one can see in table 1, the only system featuring a value higher than ours is [46]. Even though such a system could potentially lead to a higher SNR, its memory coherence time is only a few ns long, making this system unsuitable for long-distance quantum communication.
| System | Fidelity | Efficiency | In-fiber | Photon | Mean efficiency | Coherence |
| efficiency | duration | /time bin | time | |||
| Our work | 0.87 | 0.68 | 40 ns | /ns | 206 s, ms [45] | |
| Cavity-coupled | 0.97 | 0.69 | 20 s | /ns | 62 ms [44] | |
| 40Ca+ ion [43] | ||||||
| Cavity-coupled | 0.73 | 0.08 | 55 s | /ns | 200 | |
| Er3+ ion [51] | ||||||
| Cavity-coupled | 0.8 | 0.43 | 1 ns | /ns | Few ns | |
| quantum dot [46] | ||||||
| NV center [47] | 0.77 | 20 ns | /ns | 180 s [49] | ||
| Memory-coupled | 0.92 | 3 ns | /ns | 2 s | ||
| SPDC source [50] | ||||||
| DLCZ memory [48] | 40 ns | /ns | 857 s |

### v2

@@ Abstract
Our experiment starts by optically cooling and trapping individual atoms at the center of two crossed optical fiber cavities [20]. The “qubit cavity” couples the upper transition and the “herald cavity” couples the lower transition (see Fig. 1). After optically pumping the atom into state , two laser pulses induce a two-photon excitation to state , and the atom subsequently decays, emitting single photons into both cavities [31]. Both cavities have one of the mirrors with a significantly higher transmission than the other one, such that photons exit the cavities dominantly through this mirror. During the experiment, we introduce a magnetic field pointing along the qubit cavity, such that the quantization axis is defined in that direction, and single photons in a superposition of right circular and left circular polarizations are emitted into this cavity. The mirrors of the herald cavity are elliptically curved such that the cavity is birefringent with nondegenerate polarization eigenmodes [32]. The direction of the oscillating electric field of one polarization eigenmode (the mode) is oriented along the magnetic field so that this mode enhances the emission of -polarized photons. This mode is resonant to the lower transition while the other herald cavity eigenmode (the V mode) is 500 MHz blue detuned from this transition. In this situation, the two cavities enhance the emission of a single-photon polarization qubit in the qubit cavity that is entangled with the atomic spin and correlated with a -polarized single photon (herald photon) emitted in the herald cavity. The entangled state can be expressed as
@@
In order to characterize the atom-photon entangled state and the impact of the herald photon, Fig. 3 shows the density matrix of the two-qubit state for three different situations. In (a), the herald cavity is not resonant to the lower transition, and all the events with a qubit photon detection are considered. In (b), the herald cavity is resonant to the lower transition, and all the events with a qubit photon detection are considered. In (c) the herald cavity is resonant with the lower transition, too, but only the events with a photon detection in both the qubit and the herald cavities are considered. We observe how the fidelity of the atom-photon entangled state compared to an ideal Bell state increases from in (a) to in (b), and to in (c), well above the fidelity bound for entangled states () [35]. This increase is due to the enhancement of atomic decay on the lower transition, a process which is induced by the mode of the herald cavity, and which can be heralded by the detection of a polarized herald photon. The obtained heralded entangled state fidelity is limited by the decoherence of the atomic state during the atomic qubit measurement time.
@@
In conclusion, we have shown the generation of a heralded atom-photon entangled state using a single atom coupled to two optical cavities. We have observed that heralding enhances the entangled-state generation efficiency and fidelity when compared to the unheralded case. Moreover, we have used the temporal correlation information between the qubit and herald photons in order to reduce the duration of the qubit measurement window. Making use of this feature, we have observed an improvement of the entangled state fidelity from to in a noise-limited photon-qubit measurement situation, which is relevant in long-distance quantum communication.
@@ III Appendix C: Entangled state fidelity in the presence of noise
where is the fidelity of with respect to a Bell state. This equation is used to calculate the expected atom-photon entangled state fidelity in Fig. 4c in the main text.